A coloured solution is placed within a spectrometer. At 554 nm the solution shows an

absorbance of 0.69. Calculate the percentage of light that is being absorbed.


1
Expert's answer
2021-05-13T02:08:28-0400

A=-log(T/100)

0.69 = -log(T/100)

Find 10^× of both sides to cancel -log

10^(-0.69)=10^-(-log(T/100))

0.2017=T/100

Multiply both sides by 100

0.20417×100=T

T=20.417%

Percentage of light absorbed should be;

100-T = 100-20.417=79.538%